OpenStudy (anonymous):

How do I figure out "which is 5/6 of 12"?

OpenStudy (anonymous):

\[\frac{5}{6}\cdot\frac{12}{1}\]

OpenStudy (anonymous):

interpret 'of' as multiply so its 5 * 12 -- 6

OpenStudy (anonymous):

thanks!!!!!!